OpenStudy (anonymous):

x^2+7x+12 factorise completely please help me.

OpenStudy (anonymous):

(x+4)(x+3)

OpenStudy (anonymous):

how. plz show full woking

OpenStudy (anonymous):

look at the last term first

OpenStudy (anonymous):

what are the multiples of 12? (what multiply together to make 12?)

OpenStudy (anonymous):

1,12 3,4 6,2

OpenStudy (anonymous):

Do any of those numbers add or subtract from their partner to get you 7, the middle term?

OpenStudy (anonymous):

only 3 and 4

OpenStudy (anonymous):

since 12 is positive, and only a negative number times a negative number -or- a positive number times a positive number get you a positive, 3 and 4 are either -3,-4 or 3,4

OpenStudy (anonymous):

Now, is 7 positive or negative?

OpenStudy (anonymous):

it's positive

OpenStudy (anonymous):

so 3 and 4 have to be positive too. if they were negative, it would be a negative 7

OpenStudy (anonymous):

So far wh have ( 3)( 4)

OpenStudy (anonymous):

Now the last step. the first term is x^2.

OpenStudy (anonymous):

the coefficaint is 1, so your terms will be x and x. (x*x=x^2)

OpenStudy (anonymous):

well, in my point of view, it is just easier to rewrite x^2+7x+12 as x^2+4x+3x+12 Now x^2+4x+3x+12=x(x+4)+3(x+4) distributive property (x+4)(x+3)

OpenStudy (anonymous):

so finally you have (x+4)(x+3)

OpenStudy (anonymous):

p.s. once you do this a few more times, it will get a lot easier; I just wrote everything out so you wouldn't be confused by a step.

OpenStudy (anonymous):

x^2+7x+12 =x^2+4x+3x+12 =x(x+4)+3(x+4) =(x+4)(x+3)